Unit 1 Test Review

  1. Find inverse of f(x) = 3x + 4 / 5

(show step by step)

1
Expert's answer
2021-09-10T11:58:49-0400

"y=f(x)=3x+\frac{4}{5}\\
Since, \space 3x=y-\frac{4}{5}\\
x=\frac{y}{3}-\frac{4}{15}\\
Therefore, \space f^{-1}(y)=\frac{y}{3}-\frac{4}{15}"